# Beyond Dialogue: A Profile-Dialogue Alignment Framework Towards General Role-Playing Language Model

### v3

@@ Introduction
where $s$ is the prompt input, and $a$ is the corresponding label, $\odot$ represents hybrid task training, and $\pi$ denotes the trained LLMs. The total volume of training data, consistent with other comparative experiments, was regulated to be ten times that of the benchmark role-playing data, effectively managing the potential impact of data volume on LLM training. Specific training parameters are detailed in §D.
Automated Dialogue Evaluation
@@
Evaluating the effectiveness and reproducibility of role-playing is still an open problem. As discussed in §Related Work, the diversity in dialogue generation and evaluation methods complicates establishing widely accepted standards.
@@
Our another contribution is to propose an automated dialogue method, as shown on the right side of Figure 2. We first generate a role and its description, aligning it with the worldview of the role to be evaluated. Based on their profiles, we create a dialogue scenario, design the emotions, and define the relationship between the roles. Finally, the two models engage in multi-turn dialogues within this context, producing a dialogue corpus for subsequent role-playing model evaluations.
@@
During the evaluation phase, we adopted the “LLMs as judges” (Kim et al. 2023) approach, designing five tasks based on the role profile dimensions as discussed in §Related Work: Character, Style, Emotion, Relationship, and Personality. Besides, we used Human-likeness to assess whether the model’s output matches human expression and Coherence to evaluate dialogue continuity. Additionally, we introduced a role-based multiple-choice evaluation to assess the model’s role consistency across multi-turn dialogues(Lu et al. 2024). We reformatted all evaluation tasks into multiple-choice and true/false questions, clarifying criteria, streamlining the process, and reducing biases. This approach has a high level of acceptance in manual checks and closely mirrors human evaluation, making the evaluation more systematic and reliable for large-scale applications. All prompts used for dialogue generation and evaluation are available in §F.
Experiments
Experimental Setup
@@
Dataset. Following the method outlined in §Alignment Dataset Construction, we created the RP dataset by extracting 280 Chinese and 31 English roles from 123 novels and scripts, resulting in 3,552 scenario dialogues with 23,247 turns, all based on authentic sources. Compared to the other datasets listed in §A Table 7, ours has the highest number of real roles and the highest turns with real dialogues, which is significant for building a general role-playing model. Additionally, we created the CSERP dataset by deriving five alignment training tasks from each of the 3,552 sessions, resulting in five training data points per session. After aligning and adjusting role profiles with their dialogues, the data was transformed into the RPA dataset. The CC dataset included both Chinese (NaturalConv (Wang et al. 2021)) and English (DailyDialog (Li et al. 2017)) chit-chat data.
@@
Baselines. We test both open-source and proprietary advanced bilingual (Chinese-English) chatbots in our assessment framework. Our proprietary general baselines include: GPT-4o, GPT-3.5-Turbo (OpenAI 2024), Yi-Large-Turbo (Wanwu 2024), Deepseek-Chat (Deepseek 2024). These models represent the current state-of-the-art in language generation technology, covering different architectures and training methods. Additionally, we selected several baselines focused on role-playing: Index-1.9B-Character (bilibili 2024) (open-source), Baichuan-NPC-Turbo (Baichuan 2024) (proprietary), and CharacterGLM (Zhou et al. 2023) (open-source). These models are specifically designed and optimized for role-playing, capable of generating more realistic character dialogues and interactions. Furthermore, to validate the effectiveness of our framework, we trained two open-source bilingual baselines: Qwen2-7B-Instruct (Bai, Bai, and et al. 2023) and Mistral-Nemo-Instruct-2407 (Jiang et al. 2023).
@@
Metrics Design. As previously mentioned, our goal is to evaluate the model’s ability to follow the role’s profile and scenario setting, which aligns with the objectives of the derivative tasks of CSERP discussed in §A Figure 5. Consequently, the evaluation tasks and alignment tasks under the five dimensions—Character, Style, Emotion, Relationship, and Personality—are identical. Furthermore, we introduced three additional evaluation metrics: Human-likeness, Coherence, and Role Choice of the dialogues:
@@
  • Emotion and Relationship: Emotion is assessed based on Ekman’s six basic emotions (Ekman 1992), and both Emotion and Relationship are rated by GPT-4o on a scale of 0-10 based on the dialogue being evaluated. The scoring results are used to calculate the Normalized Mean Absolute Percentage Error (NMAPE) with the context labels generated by the prompt model.
@@
  • Character, Style, and Personality: Rated by GPT-4o based on the dialogue to be evaluated, referencing the candidate set of labels provided by the role profile. Notably, Personality uses the MBTI classification system and is a binary recall task, whereas Character and Style are multi-label recall tasks.
  • Human-likeness: Used to evaluate whether the role-playing model can provide a more natural and realistic interaction experience. We convert the evaluation of human-likeness into a true/false question. With few-shot prompting, GPT-4o determines whether the input dialogue is human-generated or model-generated.
  • Role Choice: Used to evaluate the recognizability of the role presentation in the dialogue. We mask the names in the dialogue and have GPT-4o select the most suitable role for the current dialogue from the four candidate roles provided.
@@
  • Coherence: Used to evaluate the logical consistency and contextual coherence of the multi-turn dialogues. GPT-4o assesses whether the dialogue is coherent based on the given context and scenario.
Main Results
@@
Model Character Style Emotion Relationship Personality Avg. ↑ Human -likeness ↑ Role Choice ↑ Coherence ↑
@@
GPT-4o 74.32 ± 1.15 81.67 ± 1.51 16.31 ± 0.48 12.13 ± 0.66 66.58 ± 4.41 78.83 ± 1.64 67.33 ± 3.95 87.33 ± 3.86 99.67 ± 0.33
GPT-3.5-Turbo 72.26 ± 1.27 73.66 ± 1.73 17.79 ± 0.56 14.17 ± 0.73 66.92 ± 4.85 76.18 ± 1.83 $3$ 4.43 83.00 ± 4.68 97.33 ± 1.17
Yi-Large-Turbo 75.13 ± 1.22 79.18 ± 1.58 16.44 ± 0.49 13.48 ± 0.67 68.25 ± 4.61 78.53 ± 1.72 47.00 ± 4.60 84.33 ± 3.67 92.67 ± 2.39
Deepseek-Chat 75.46 ± 1.14 81.49 ± 1.51 15.92 ± 0.46 12.42 ± 0.63 67.92 ± 4.57 79.30 ± 1.66 52.33 ± 4.95 83.00 ± 4.68 96.67 ± 1.00
@@
Index-1.9B-Character 73.33 ± 1.32 76.48 ± 1.50 17.99 ± 0.53 13.58 ± 0.71 66.33 ± 4.57 76.92 ± 1.73 21.67 ± 3.96 78.67 ± 5.14 69.67 ± 3.85
CharacterGLM-6B 73.36 ± 1.28 76.08 ± 1.55 18.58 ± 0.55 14.27 ± 0.79 67.33 ± 4.34 76.79 ± 1.70 16.00 ± 2.38 81.00 ± 4.40 25.67 ± 3.48
Baichuan-NPC-Turbo 75.19 ± 1.23 79.15 ± 1.38 17.24 ± 0.51 13.10 ± 0.69 65.33 ± 4.84 77.87 ± 1.73 56.00 ± 4.66 86.33 ± 4.90 99.00 ± 0.56
@@
Mistral-Nemo-Instruct-2407 74.12 ± 1.17 77.04 ± 1.48 17.00 ± 0.43 13.50 ± 0.67 67.00 ± 4.30 77.53 ± 1.61 53.67 ± 4.66 82.67 ± 4.77 74.33 ± 3.77
+ RP & CC 71.56 ± 1.26 74.66 ± 1.51 17.36 ± 0.49 12.58 ± 0.69 68.17 ± 4.32 76.89 ± 1.65 53.67 ± 3.76 86.33 ± 4.22 86.33 ± 2.42
+ RPA & CC 74.44 ± 1.14 77.63 ± 1.40 16.74 ± 0.46 12.07 ± 0.67 69.50 ± 4.31 78.55 ± 1.60 56.00 ± 3.41 85.00 ± 4.49 91.67 ± 1.80
+ RPA & CC & CSERP 74.58 ± 1.28 78.47 ± 1.45 16.62 ± 0.48 11.38 ± 0.67* 69.08 ± 4.46 78.83 ± 1.67 59.00 ± 4.46 87.00 ± 4.73 92.67 ± 1.59
@@
+ RP & CC 74.91 ± 1.21 78.59 ± 1.39 16.52 ± 0.48 12.28 ± 0.67 67.00 ± 4.31 78.34 ± 1.61 56.67 ± 3.85 82.00 ± 4.90 90.00 ± 1.59
+ RPA & CC 76.43 ± 1.18 81.28 ± 1.37 16.10 ± 0.45 12.35 ± 0.67 67.83 ± 4.36 79.42 ± 1.61 62.33 ± 3.21 87.33 ± 3.77 92.67 ± 1.85
+ RPA & CC & CSERP 78.67 ± 1.12* 82.52 ± 1.33* 15.68 ± 0.5* 11.22 ± 0.72* 69.67 ± 4.27 80.79 ± 1.59* 64.33 ± 3.80* 87.33 ± 3.74 99.00 ± 0.56
Table 2: Main results of Beyond Dialogue. We report the average scores with their standard error of the mean (SEM). Bold numbers indicate the highest scores, while underlined numbers are the best in the group, * means the improvements over the untrained baseline are statistically significant (achieving $p<0.05$ in t-test). The Avg. column represents the mean of the Character, Style, Emotion, Relationship, and Personality metrics, which are directly defined within the profile, with Emotion and Relationship calculated as 1-NMAPE. All dialogues are evaluated by GPT-4o as judges.
We conducted a total of 300 bilingual evaluations for each model, with each evaluation consisting of a scenario containing five turns of dialogue. Specific evaluating parameters are detailed in §D. The evaluation process was fully automated, with GPT-4o generating chat roles and scenarios, and engaging in multiple turns of dialogue with the evaluated model. The generated dialogues were input into an automatic evaluation pipeline to obtain quantitative performance metrics for the role-playing models. These evaluation results were used to calculate the average performance of each model across five dimensions and three additional metrics.
@@
In our experiments, the model trained with unaligned role-playing data (+RP & CC) exhibited training biases, leading to limited or even diminished performance in the unaligned dimensions of Character, Style, and Personality. However, the model trained with aligned data (+RPA & CC) showed significant improvements in the character and style dimensions, with average scores surpassing both general-purpose and expertise baselines, indicating the importance of aligned data in mitigating training bias.
@@
Overall, these results emphasize the importance of aligned dialogues and role profiles in role-playing training and demonstrate the excellent performance of the “Beyond Dialogue” training tasks in general role-playing dimensions.
Further baseline results can be found in the supplementary tables provided in §B. Case studies are available in §G. For a detailed account of the manual verification process used in GPT-4o evaluations, please refer to §C.
Alignment Results
To explore the model’s capability in fine-grained alignment for dialogue and role profiles, we randomly sampled 100 data points from each dimension of CSERP in the evaluation results (§Main Results). Using GPT-4o’s results as the reference labels, we present the alignment results of the comparable baseline model and two improved models in Table 3. The alignment for Character, Style, and Personality was assessed using recall, while Emotion and Relation were evaluated using NMAPE.
@@
The fine-tuned Mistral-Nemo model, using the full dataset, achieved top performance, closely matching GPT-4o’s 89.4% effectiveness. Both Mistral and Qwen showed notable improvement, with recall rates for Character and Style increasing by over 30%, while NMAPE for Emotion and Relation decreased significantly. This enhancement allowed the 7B-parameter Qwen2 model to surpass the 9B-parameter Yi1.5 and GLM4 models in alignment tasks.
However, when trained on the RPA & CC or RP & CC datasets alone, both Qwen and Mistral struggled with alignment tasks. Their reasoning capabilities declined, resulting in ineffective outcomes. Detailed case studies are in §G.
@@
Ablation Study
To assess the contribution of the five CSERP alignment tasks within the Beyond Dialogue framework, we conducted an ablation study. Each task was removed individually, with its training data volume replaced by an equivalent amount of CC data to maintain consistency. Subsequently, we evaluated the model across the five role-playing dimensions of CSERP, using the same evaluation methods as in Table 2.
Table 4 presents the results of the ablation experiments for CSERP training tasks. Notably, the ablation of w/o Char. and w/o Styl. have the most significant impact, with a marked decline in the performance of Character and Style, respectively. The other three metrics also showed varying degrees of decline, indicating that these fine-grained alignment tasks are crucial for the model’s general role-playing capability.
(a) Average Score

### v6

@@ 3.2 Supervised Finetuning
where $s$ is the prompt input, and $a$ is the corresponding label, $\odot$ denotes training dataset mixing, and $\pi$ denotes the trained LLMs. The total volume of training data, consistent with other comparative experiments, was regulated to be ten times that of the benchmark role-playing data, effectively managing the potential impact of data volume on LLM training. Training parameters are detailed in § D.
3.3 Automated Dialogue Evaluation
@@
Evaluating the effectiveness and reproducibility of role-playing is still an open problem. As discussed in § 2, the diversity in dialogue generation and evaluation methods complicates establishing widely accepted standards.
@@
Our next contribution is to propose an automated dialogue method, as shown on the right side of Figure 3. We first generate a role and its description, aligning it with the worldview of the role to be evaluated. Based on their profiles, we create a dialogue scenario, design the emotions, and define the relationship between the roles. Finally, the two models engage in multi-turn dialogues within this context, producing a dialogue corpus for subsequent role-playing model evaluations. One of the models in the dialogue is GPT-4o, which compared to human evaluators, provides a low-cost and standardized approach, ensuring consistent evaluation across models (Wang et al., 2024b; Lu et al., 2024).
@@
During the evaluation phase, we adopted the “LLMs as Judges” (Kim et al., 2023) approach, designing five tasks based on the role profile dimensions as discussed in § 2: Character, Style, Emotion, Relationship, and Personality. Besides, we used Human-likeness to assess whether the model’s output matches human expression and Coherence to evaluate dialogue continuity. Additionally, we introduced a role-based multiple-choice evaluation to assess the model’s role consistency across multi-turn dialogues.We reformatted all evaluation tasks into objective questions, yielding lower variance and closer alignment with human assessments, compared to subjective ones (as shown in § B.1). This approach has a high level of acceptance in manual checks and closely mirrors human evaluation, making the evaluation more systematic and reliable for large-scale applications. All prompts used for dialogue generation and evaluation are available in § F.3.
4 Experiments
4.1 Experimental Setup
@@
Dataset. Following the method in § 3.1, we created the role-playing dialogue dataset (RP) by extracting 280 Chinese and 31 English roles from 123 novels and scripts, resulting in 3,552 scenario dialogues with 23,247 turns, all from authentic sources. Compared to the other datasets listed in § A.3 Table 7, ours has the highest number of real roles and the highest turns with real dialogues, which is significant for building a general role-playing model. Additionally, we created the sentence-level alignment dataset (CSERP) by deriving five alignment training tasks from each of the 3,552 sessions, resulting in five training data points per session. After alignment, only 4.2% of the dialogues in the of the dialogues in the Character, Style, and Personality dimensions were fully consistent with the predefined role profiles (see § A.3 for dataset statistics). Once the role profiles were aligned and adjusted with their dialogues, the data was transformed into the aligned role-playing dialogue dataset (RPA). The chit-chat dataset (CC) included both Chinese (NaturalConv (Wang et al., 2021)) and English (DailyDialog (Li et al., 2017)).
@@
Baselines. We test both open-source and proprietary advanced bilingual (Chinese-English) chatbots in our assessment framework. Our proprietary general baselines include: GPT-4o, GPT-3.5-Turbo (Hurst et al., 2024), Yi-Large-Turbo (Young et al., 2024), Deepseek-Chat (Bi et al., 2024). These models represent the current state-of-the-art in language generation technology, covering different architectures and training methods. Additionally, we selected several baselines focused on role-playing: Index-1.9B-Character (bilibili, 2024) (open-source), Baichuan-NPC-Turbo (Yang et al., 2023) (proprietary), and CharacterGLM (Zhou et al., 2024) (open-source). These models are specifically designed and optimized for role-playing, capable of generating more realistic character dialogues and interactions. Furthermore, to validate the effectiveness of our framework, we trained two open-source bilingual baselines: Qwen2-7B-Instruct (Yang et al., 2024) and Mistral-Nemo-Instruct-2407 (MistralAI, 2024).
@@
Metrics Design. As previously mentioned, our goal is to evaluate the model’s ability to follow the role’s profile and scenario setting, which aligns with the objectives of the derivative tasks of CSERP discussed in § A.1 Figure 5. Consequently, the evaluation tasks and alignment tasks under the five dimensions — Character, Style, Emotion, Relationship, and Personality — are identical. Furthermore, we introduce four additional evaluation metrics: Human-likeness, Coherence, Role Choice, and Win-Rate. These metrics are evaluated from the perspective of the entire session’s dialogue. Notably, while Win-Rate is assessed manually, the remaining metrics are automatically evaluated using LLMs:
@@
  • Emotion and Relationship: Both Emotion, based on Ekman’s six basic emotions (Ekman, 1992), and Relationship are rated by LLM on a 0-10 scale based on the evaluated dialogue. The scores are then used to calculate the Normalized Mean Absolute Percentage Error (NMAPE) against context labels generated by the prompt model.
@@
  • Character, Style, and Personality: Rated by LLM based on the dialogue and the role profile’s candidate labels. Personality is assessed using the MBTI classification (binary recall), while Character and Style are multi-label recall tasks.
  • Human-likeness: Evaluates the naturalness and realism of the interaction. LLM determines whether the dialogue is human or model-generated using few-shot prompting.
  • Role Choice: Assesses how well the role is recognized in the dialogue. LLM selects the most appropriate role from four candidates after masking the role names in the dialogue.
@@
  • Coherence: Evaluates the logical consistency and contextual coherence of multi-turn dialogues. LLM checks whether the dialogue is coherent within the given context.
  • Win-Rate: Assesses the model’s performance through pairwise human comparisons with GPT-4o, where multiple annotators cast votes on their preference, and the final decision is based on the majority agreement.
4.2 Main Results
@@
Model Character Style Emotion Relationship Personality Qualification -Rate Human -likeness ↑ Role Choice ↑ Coherence ↑ Win-Rate ↑ (vs. GPT-4o)
@@
GPT-4o 74.32 ± 1.15 81.67 ± 1.51 16.31 ± 0.48 12.13 ± 0.66 66.58 ± 4.41 46.33 ± 2.88 67.33 ± 3.95 87.33 ± 3.86 99.67 ± 0.33 N/A
GPT-3.5-Turbo 72.26 ± 1.27 73.66 ± 1.73 17.79 ± 0.56 14.17 ± 0.73 66.92 ± 4.85 39.0 ± 2.82 $3$ 4.43 83.00 ± 4.68 97.33 ± 1.17 37.00 ± 2.79
Yi-Large-Turbo 75.13 ± 1.22 79.18 ± 1.58 16.44 ± 0.49 13.48 ± 0.67 68.25 ± 4.61 49.0 ± 2.89 47.00 ± 4.60 84.33 ± 3.67 92.67 ± 2.39 45.67 ± 2.88
Deepseek-Chat 75.46 ± 1.14 81.49 ± 1.51 15.92 ± 0.46 12.42 ± 0.63 67.92 ± 4.57 49.33 ± 2.89 52.33 ± 4.95 83.00 ± 4.68 96.67 ± 1.00 49.67 ± 2.89
@@
Index-1.9B-Character 73.33 ± 1.32 76.48 ± 1.50 17.99 ± 0.53 13.58 ± 0.71 66.33 ± 4.57 41.67 ± 2.85 21.67 ± 3.96 78.67 ± 5.14 69.67 ± 3.85 44.33 ± 2.87
CharacterGLM-6B 73.36 ± 1.28 76.08 ± 1.55 18.58 ± 0.55 14.27 ± 0.79 67.33 ± 4.34 36.0 ± 2.77 16.00 ± 2.38 81.00 ± 4.40 25.67 ± 3.48 37.00 ± 2.79
Baichuan-NPC-Turbo 75.19 ± 1.23 79.15 ± 1.38 17.24 ± 0.51 13.10 ± 0.69 65.33 ± 4.84 49.33 ± 2.82 56.00 ± 4.66 86.33 ± 4.90 99.00 ± 0.56 65.00 ± 2.75
@@
Mistral-Nemo-Instruct-2407 74.12 ± 1.17 77.04 ± 1.48 17.00 ± 0.43 13.50 ± 0.67 67.00 ± 4.30 39.0 ± 2.82 53.67 ± 4.66 82.67 ± 4.77 74.33 ± 3.77 34.33 ± 2.74
+ RP & CC 71.56 ± 1.26 74.66 ± 1.51 17.36 ± 0.49 12.58 ± 0.69 68.17 ± 4.32 34.67 ± 2.75 53.67 ± 3.76 86.33 ± 4.22 86.33 ± 2.42 62.33 ± 2.80
+ RPA & CC 74.44 ± 1.14 77.63 ± 1.40 16.74 ± 0.46 12.07 ± 0.67 69.50 ± 4.31 43.67 ± 2.86 56.00 ± 3.41 85.00 ± 4.49 91.67 ± 1.80 65.00 ± 2.75
+ RPA & CC & CSERP 74.58 ± 1.28 78.47 ± 1.45 16.62 ± 0.48 11.38 ± 0.67* 69.08 ± 4.46 47.33 ± 2.88* 59.00 ± 4.46 87.00 ± 4.73 92.67 ± 1.59 69.33 ± 2.66
@@
+ RP & CC 74.91 ± 1.21 78.59 ± 1.39 16.52 ± 0.48 12.28 ± 0.67 67.00 ± 4.31 38.0 ± 2.8 56.67 ± 3.85 82.00 ± 4.90 90.00 ± 1.59 64.00 ± 2.77
+ RPA & CC 76.43 ± 1.18 81.28 ± 1.37 16.10 ± 0.45 12.35 ± 0.67 67.83 ± 4.36 44.67 ± 2.87 62.33 ± 3.21 87.33 ± 3.77 92.67 ± 1.85 67.33 ± 2.71
+ RPA & CC & CSERP 78.67 ± 1.12* 82.52 ± 1.33* 15.68 ± 0.5* 11.22 ± 0.72* 69.67 ± 4.27 56.33 ± 2.86* 64.33 ± 3.80* 87.33 ± 3.74 99.00 ± 0.56 71.00 ± 2.62
Table 2: Main results of Beyond Dialogue. We report the average scores with their standard error of the mean (SEM). Bold numbers indicate the highest scores, while underlined numbers are the best in the group, * denotes statistically significant improvements over the untrained baseline (achieving $p<0.05$ in t-test). The Qualification Rate (QR) indicates the proportion of role-playing dialogues that align with the predefined profile, scoring above 60 in each C-S-E-R-P dimension, with E and R calculated as 1-NMAPE. A darker background indicates better performance. More baseline results are provided in § B.2 Table 10. All dialogues are evaluated by GPT-4o as judges, with supplementary evaluations provided by Claude-3-Opus in § B.2 Table 11, yielding comparable results.
We conducted 300 independent bilingual evaluations for each model, with each evaluation consisting of a scenario containing five turns of dialogue, all involving new roles and scenarios. Specific evaluating parameters are detailed in § D. The evaluation process was fully automated, with GPT-4o generating new chat roles and scenarios, and engaging in multiple turns of dialogue with the evaluated model (as illustrated by role-playing dialogue cases in § G.1). The generated dialogues were input into an automatic evaluation pipeline to obtain quantitative performance metrics for the role-playing models. These evaluation results were used to calculate the qualification rate of each model across five dimensions and four additional metrics.
@@
In our experiments, the model trained with unaligned role-playing data (+RP & CC) exhibited training biases, resulting in limited or even diminished performance in Qualification-Rate. In contrast, the model trained with aligned data (+RPA & CC) achieved significant improvements, emphasizing the importance of data alignment in mitigating training bias. Furthermore, the model trained with additional aligned data (+RPA & CC & CSERP) outperformed the original model by 19%, and surpassed GPT-4o by 10%.
@@
The inclusion of Win-Rate metrics from pairwise human comparisons against GPT-4o offers a more comprehensive evaluation of model performance. Models fine-tuned with our framework demonstrate substantial improvements in human-likeness, with significantly higher win rates compared to their baseline performance. According to annotator feedback, models trained with our framework exhibit more natural and human-like conversational styles, in contrast to the “mechanical” responses typically seen in general-purpose models.
These qualitative improvements are further supported by manual verification, which demonstrates strong inter-rater agreement between GPT-4o and human evaluations (§ C.2, Table 12), thereby confirming the reliability of our approach.
Overall, these results highlight the importance of aligning role profiles and dialogues in role-playing training, and demonstrate that “beyond dialogue” tasks help models better adhere to role profiles.
4.3 Alignment Results
To explore the model’s capability in fine-grained alignment for dialogue and role profiles, we randomly sampled 100 data points from each dimension of CSERP in the evaluation results (§ 4.2). Using GPT-4o’s results as the reference labels, we present the alignment results of the comparable baseline model and two improved models in Table 3. The alignment for Character, Style, and Personality was assessed using recall, while Emotion and Relation were evaluated using NMAPE.
The fine-tuned Mistral-Nemo model, achieved top performance, closely matching GPT-4o’s 89.4% effectiveness. Both Mistral and Qwen showed notable improvement, with recall rates for Character and Style increasing by over 30%, while NMAPE for Emotion and Relation decreased significantly. This boost allowed the 7B-parameter Qwen2 model to surpass the 9B-parameter Yi1.5 and GLM4 models in alignment tasks.
@@ 4.3 Alignment Results
However, when trained on the RPA & CC or RP & CC datasets alone, both Qwen and Mistral struggled with alignment tasks. Their reasoning capabilities declined, resulting in ineffective outcomes. Detailed case studies are in § G.2.
@@
4.4 Ablation Study
To assess the contribution of the five CSERP alignment tasks within the Beyond Dialogue framework, we conducted an ablation study. Each alignment task in $D_{a}$ was removed individually, with its training data volume replaced by an equivalent amount of $D_{c}$ to maintain consistency. We evaluated the model on the five role-playing dimensions using the same methods in Table 2.
Table 4 presents the results of the ablation experiments for CSERP training tasks. Notably, the ablation of w/o Char. and w/o Pers. have the most significant impact, with a marked decline in the performance of Qualification Rate. The other three metrics also showed varying degrees of decline, indicating that these fine-grained alignment tasks are crucial for the model’s general role-playing capability.
(a) Qualification Rate